"Configuration Mgmt Template"
IT Services Weekly Project Status Report 13-Feb-06 Hot Issues Initia Project Project % (Must be completed for Projects Project Description Funding tive Sponsor Manager Status Project Start Project Finish Complete Status Yellow or Red status) Active Directory Kit Develop a “Self Serve” kit for clients to -- -- Barry M Caren K Yellow 5-Dec-05 TBD 5% 2/10: Will be working with the new WST manager 2/10: Sponsor changed to Barry migrate into Active Directory. for John F Barry Magsanay. for interim. Resources still questionable. Active Directory Assist specific clients with migration to -- -- Barry M Caren K Green 1-Sep-05 31-Aug-06 30% 2/10: Tim Posten met with CRC. Finalizing Migrations Active Directory. for John F project. AMCOM Upgrade AMCOM paging system. -- -- Jay K Michelle C Green 15-Nov-05 31-Mar-06 70% 2/10: Substantial progress made in all project areas this week: 1) Communication and training plan agreed upon with Hospital for OnCall admin staff.Devlopment for both in progress. 2)Tesing underway for web-interface and operators. 3) Training negotiated with AMCOm for operators, scheduled to start week of 27th. 4) On-Target for scheduled app cut-over. Audit Response - Remediate security audit findings -- 31 Bill C Caren K Green 5-Aug-06 31-Aug-06 80% 2/10: Response from Bruce V and Steven. Internal Making prgress. Audit Response - Coordinate IT Services responses for -- 32 Jay K Caren K Green 5-Aug-06 31-Aug-06 80% 2/10: Change in Sponsor to Jay Kohn External (PWC) recommendations from PWC. Change Pilot new Change Management -- 18 Jay K Bill H Complete 1-Mar-05 31-Dec-05 100% Done. Pilot live 12/14/05. Management Phase process and system to improve IT 1 Services’ ability to track, approve, and implement changes to technologies, architecture and service levels. Pilot group will be Shared Application Systems. Change Implement additional requirements for -- 19 Jay K Bill H Green 1-Jan-06 10-Mar-06 82% 2/10: Received bug fix release from Infra on 2/6. 2/10: Status changed from Management Phase Change Management tools and Regression testing underway. Quality of release Yellow to Green since release 2 processes. Complete phased rollout so far looks good. Training schedule for Go Live looks good to IT Services and Administrative has been published. Systems. 2/6: Seven new participants have joined the pilot (from the Telecom and Unix teams). Training was completed for them this week. The next Infra release is arriving on 2/6. 1/30: no update since last week 1/23: Work is continuing with the vendor on issues to be resolved in a release due on 2/6. A second wave of users will join the pilot on 1/27 - they will be trained this week. A go-live is planned for March. 1/7/2011 page 1 of 7 IT Services Weekly Project Status Report 13-Feb-06 Hot Issues Initia Project Project % (Must be completed for Projects Project Description Funding tive Sponsor Manager Status Project Start Project Finish Complete Status Yellow or Red status) Continuing Studies - Transition Continuing Studies Web & Client -- John F Caren K Red 1-Feb-06 1-Mar-06 0% 2/10: Project timeline slipping due to an undefined 2/10: No WST resources have eCommerce database server from a 1 tier windows WST resource. been defined. Following up with Implementation server environment to a more secure, John F. PCI compliant 2 tier windows server environment with firewalls. Desktop Backup Transition PC desktop to third-party -- 22 Jan C Caren K Green 1-Apr-05 31-Dec-06 50% 2/10: Change in sponsor, will now be Jan C. vendor for desktop backup. Update Mac desktop backup on TSM. Desktop CRC desktop configuration for all of SGG 14 Jan C Steve L Complete 1-Mar-05 1-Dec-05 100% Done. Configuration Mgmt CRC clients. Desktop Support the application and SGG 14 Jan C Steve L Green 1-Jan-06 1-Mar-06 50% 2/10: MedIRT review complete, testing and pilot in 2/10: Barry Magsanay named Configuration Mgmt- maintenance of Category A (HIPAA) MedIRT in 3/06, Web site and documentation first as new sponsor and owner -- will Phase 2 required security settings, as well as drafts complet work on coordination noted allow local customization of the DCM 2/3: MedIRT integration XML document complete. below profiles, allowing local IT groups to Getting Started and Service Web Site under 2/3: Who will own this service tailor the profile to their needs. construction and coordiante the integration with other Desktop services such as BigFix and Host Reg Disaster - Review, prioritize, and fulfill highest Jung 48 John F Steve L Green 1-Sep-05 1-Jun-06 40% 2/10: Audit letter favorably recieived vis a vis 2/10: JohnF and SteveJ to meet Recovery/Business priority EOC data needs EOC progress. Final 3 candidates to be referred to to handoff to new sponsor in ITS. Continuity - Determine disaster recovery SteveJ and Keith Perry Project charter review indicated strategies for information deemed to 2/3: Hiring process underway for FTE. Data Matrix that 2nd half of the work (50%) be of highest priority to the for EOC is completed, and all EOC member will begin with new hire University’s emergency response meetings held 2/3: Need to work on getting - Implement one or more business data to EOC, possible continuity services partnerships assistance to EHandS for File - Continue detailed planning for Server disaster recovery and business continuity EMC Storage Resolve all outstanding storage -- -- John F Steve L Red tbd NA NA 2/10: New DMX mitigation using the Clariion is 2/10: Resource risk still very issues, including Forsythe SAN underway. Further refining of the "no dynamic high, but interviews and audits Replacement (Migration of Forsythe spare" and new disk for new BCV options. 600B to are helping to define a direction. SAN to Cisco equipment dual- D migration progress slowed due to DMX effort Project still needs better form for attaching each host to eliminate single and operational priorities scope, tasks, dates, priorities. point of failure) and EMC Storage 2/3: AS DMX priorities identified and action plan Data still on 600B Strategy. underway. ATA Disk 600B - 600D migration 2/3: Fourth week as RED on underway. Several concrete interviews for extra Client Relations Weekly Client resources Report. Staffing and scope risks still exist. Still difficult to come up with comprehensive plan, hard to tease out operational issues from project issues Firewall Operations Project management of new clients to -- -- Jay K Michelle C Green 30-Oct-05 NA NA 2/10: 3-Feb Management - the firewall service, additions of zones 1) Substantial follow-up with PeopleSoft staff on Client & Other to existing firewall projects and other project to Load Balance 2 servers - concern about Projects client firewall initiatives. processes. 2) Weekly meeting with Liz G on upcoming client projects. 1/7/2011 page 2 of 7 IT Services Weekly Project Status Report 13-Feb-06 Hot Issues Initia Project Project % (Must be completed for Projects Project Description Funding tive Sponsor Manager Status Project Start Project Finish Complete Status Yellow or Red status) Firewall Operations A series of discrete projects focussed -- -- Jay K Michelle C Green 30-Oct-05 NA NA 2/10: Continue editing and revision of FW service Management - on the improvement of the operational description based on feedback from team. Team Operations Process processes associated with the firewall will follow-up with Eric N on refining processes Development service. This includes rates where security is involved. development and associated tasks 2) Drafted communication for clients around the such as developing tools to automate proposed change in "template" pocedures. the metrics capturing capabilities for 3) Mark M and I met with Nan M to get an biling purposes on the firewall, overview of PAC and its applicability for Firewall upgrading existing client facing service. documentation and processes, etc. Firewall Operations Develop process for capturing, -- -- Jay K Michelle C Green 30-Oct-05 NA NA 2/10: Verification of PTAs continues with clients Management - Rate formalizing, and rolling out rates for for billing. Development firewall service. HIPAA Secure Pilot secure email service that allows Provost 30 John F Steve L Green 1-Jan-05 1-Jun-06 70% 2/10: Performance issues still need to solved at 2/10 All features in place. Need Email Stanford users to send encrypted the email filter level. Meeting with Sponsors to do to work on "self identification" for email and mitigate HIPAA requirement first status check of 2006 next testing round to facilitate encrypted email. 2/3: UAT in progress, troubleshooting will begin for 2/3: Med School reqest for Hi performance issues. New Testing/Development Volume group programming still Voltage appliance purchase approved needs to be done. Hardware Host Self Implement automated method to verify SGG 29 Jan C Steve L Green 1-Sep-05 1-Jun-06 20% 2/3: Need to develop milestones for code 2/3: Significant dependencies on Registration machines are not at risk prior to 2/3: Health Check tool complete NetDB and Network work still joining Stanford network. Ports exist existing implementation to java. Inventory Database Inventory Database -- -- Jay K Bernadette D Green 1-Sep-05 31-Aug-06 55% 2/10: Data reconciliation issues resulted in additional changes requiring another release before go live. Doc/Dictionary is complete except for integration of mentioned additional changes. 2/2: New release implemented. Testing to start. Documentation and Dictionary draft will be complete and ready for review by 2/10. Kerberos Migration Retire Stanford's Kerberos 4 cell and SGG 33 John F Jon P Green 1-Jan-06 1-Apr-08 20% 2/10: Released new S/Ident and kftgt packages, (4 --> 5) move all centrally-provided services released webauth 3.4.1 (bug fix release) 2/6: that rely on Kerberos 4 for Designing work around for bad behavior for authentication to use Kerberos 5. SPNEGO/Windows/MSIE. Brad reviewing UI for OpenAFS extensions and Weblogin interface tweaks. Loadbalancing Global Load Balancing Service for -- 21 John F Bernadette D Yellow 1-Jan-06 31-Aug-06 5% 2/10: Project kick-off meeting resulted in additional '2/10: Has been red on client Service Clients: issues with project charter. Target is to resolve report for a number of weeks. It Initiate load balancing service to issues with project sponsor by 2/17. should remain in Yellow until enhance system performance and 2/2: Draft project charter reviewed with project project charter is approved and availability sponsors. Collecting expense info are basis for timeline is set. determining LB rates. Outsource Goal: work with clients to identify -- -- John F Michelle C Green 1-Oct-05 31-May-06 80% 2/10: Meeting set with procurement for negotiating Highspeed Printing suitable alternatives to Forsythe balnket agreements with vendors. printing, assist clients with the 2) Confirmed with vendors need to use PGP for transition to the new method, and securing data. shutdown the forsythe print service by 3 )Conference call with ALC vendor to confirm 6/1/06. tenative schedule for testing. 4) Will communicate plan this week with clients. 1/7/2011 page 3 of 7 IT Services Weekly Project Status Report 13-Feb-06 Hot Issues Initia Project Project % (Must be completed for Projects Project Description Funding tive Sponsor Manager Status Project Start Project Finish Complete Status Yellow or Red status) Pathworks -- -- Jay K TBD Green 1-Mar-06 1-May-06 0% 2/6: Need to assign PM 2/13: Verifying start date in Implementation March PMO BPR Develop a set of processes and tools -- 7 Jan C Joyce D Green 5-Oct-05 6-Aug-06 50% 2/13: Project Financials received, under review. to more effectively manage projects in Contract PM to start 2/15 for implementation IT Services. 2/6: Financial tracking process clarified - expect project finance reports by mid-Feb. RWC - Data Center Prepare Data Center in Redwood City -- 44 Jay K Bob M Yellow 1-Oct-05 30-Dec-06 20% 2/6: Design team to meet on 2/7 to begin next 1/30: Pending Board Approval in Buildout (RWC) for move (get the building round of data center design. mid-February. 12/30 construction ready) date assumes no delays with Board and/or fund approval RWC - Strategy Coordinate move of services and -- 46 John F Jon P Green 1-Sep-05 1-Jun-07 5% 2/10: Steering Committee met; general agreement 1/23: Need signoff from Bill, Bob equipment to Redwood City. to reduced scope of service move. Need to circle & Randy on charter. back and confirm reduced budget estimates with service groups. Revised timetable—data center not available until Spring 07 2/6 Rebudgeted services move for reduced scope. Bill meeting with Randy on 2/6 to discuss. Shibboleth The project will deploy a Shibboleth SGG 34 John F Jon P Green 1-Jan-06 31-Aug-06 5% 2/10: Met to discuss Admin Guide changes for 1/30: Second week as YELLOW infrastructure that will support the Shib—Bruce working on a proposal. Packaging on Client Relations Weekly authentication of Stanford users to work to start this week. Met w/Lois. Client Report access authorized online web 2/6: Shib2 rebuilt and shib software installed. 1/'23: PM assigned to Jon P resources at other participating InQueue (test) federation joined and configured on institutions. The project will also our servers. Bruce to give Shib presentation 2/16. provide documentation for Stanford Off of client support yellow status. service providers who want to make their web applications available to users across federations. Sybase Elimination Complete the port of NetDB and SGG 15 John F Jon P Green 1-Jan-05 1-Jun-06 65% 2/10: Transitioning release process from Dmitri to 1/23: CRC to provide testing for NetDB related technologies from Sybase to Joel. First alpha a little rocky, but released and resource in Feb Oracle, and refactor the database quality improving. Replacements for mother and 1/6: additional testing resources interface layer of the NetDB system. tachyon build. 2/6: First QA label to be pressed required. Monday. Network screen not yet functional, all others are. 1/7/2011 page 4 of 7 IT Services Weekly Project Status Report 13-Feb-06 Hot Issues Initia Project Project % (Must be completed for Projects Project Description Funding tive Sponsor Manager Status Project Start Project Finish Complete Status Yellow or Red status) TimeTracking a. Standardize, reconfigure -- 43 Sam S Bernadette D Yellow 1-Jan-06 1-Jun-06 10% 2/10: No update. 1/25: Project Manager assigned Phase 2 Timetracking application 2/2: Steering Committee scheduled for 2/10. to Bernadette b. Ensure current categories and Transition of documents from Caren K. to 1/20: SAS person has not been configurations are useful to staff as Bernadette complete. hired. Sent out an e-mail will a well as rate setting proposed plan. Jan will speak to c. Educate / motivate managers to use John F directly about a solution. application for staff resource management and future planning Unified messaging Pilot to determine replacement of -- -- Jay K Michelle C Yellow ?? ?? 0% 2/10:Project will be delayed until April-May WSG staff are critical for this existing OCTEL Voice Mail timeframe as necessary resources are not project . Due to their staff messaging system. available at this time. shortage and existing commitments, they can not assist until April-May. This will not affect funding, and a delay at this time is preferable to starting with resources not fully committed. Status is defined as the Percentage Complete, and then Green, Yellow or Red: Complete The project has been closed and is 100% complete. Hold The project is postponed or cancelled. No work is being done at the time. Green The project is on track to be completed as scheduled and within scope and budget. Confidence level is 90% or greater. Yellow The project is at risk. It should still hit its completion targets, but there are known problems that threaten to take it off course. Confidence level is 70-90%. Red The project is off track. There have already been problems impacting completion or they can been seen on the horizon. Confidence level is below 70%. Descriptions of the individual projects can be found at: www.stanford.edu/dept/its/organization/projectmanagement To receive this report weekly from the IT Services Project Management Office, subscribe to: email@example.com 1/7/2011 page 5 of 7 ResComp Jan C Michelle C TBD TBD 0% Green Security Interventio n and Remediati on system RWC - Develop a -- 45 Jay K Jon P Hold 1-Sep-05 1-Jun-07 Staff tentative Relocation plan for the move of IT Services employee s to Redwood City. 1/23 :Not yet a project, probably will result in an initial meeting, and some consolidati on of existing web based info. 5% 1/30: On hold until staff move strategy defined. Staff moves related to Data Center move incorporat ed in RWC- Services Move project.